OpenStudy (anonymous):

Find the area of a parallelogram if a base and corresponding altitude have the indicated lengths. Write your answer as a mixed number. Base 3 1/2 feet, 3/4 altitude feet. PLEASE HELPP!!!!!! THANKS :)

OpenStudy (ash2326):

\[Area\ of\ parallelogram\ = Base\ \times\ Altitude\] \[Base= 3\frac{1}{2}=3+\frac{1}{2}=\frac{3\times 2}{1\times 2}+\frac{1}{2}=\frac{6}{2}+\frac12=\frac72\] \[Altitude=\frac{3}{4}\] so \[Area=\frac72 \times \frac34\] I think you can solve now, can't you?

OpenStudy (anonymous):

just multiply the fractions right?

OpenStudy (ash2326):

Yeah, that's it:)

OpenStudy (anonymous):

Yes.

OpenStudy (ash2326):

Don't forget the units, feet^2

OpenStudy (anonymous):

Sorry, I thought ash left :P

OpenStudy (anonymous):

so is it 4.6?

OpenStudy (ash2326):

\[21/8 \ne\ 4.6\] Check again!!

OpenStudy (anonymous):

2.625 sqft?

OpenStudy (ash2326):

Yeah:D great work:D

OpenStudy (anonymous):

thank you sooo much :D

OpenStudy (anonymous):

its saying that the answer is wrong

OpenStudy (ash2326):

Type the answer in mixed fraction or fraction form \[\frac{21}{8}=2\frac{5}{8}\]

OpenStudy (anonymous):

yes got it! @ash2326

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Thank you sooo much!!!